# A new long gamma-ray burst formation pathway at solar metallicity

### v1

@@ 1 Introduction
Due to their highly-beamed and energetic nature, LGRBs have been observed up to $z\sim 9.4$ (Cucchiara et al., 2011) and have been used to estimate the star formation rate at high redshifts. However, the LGRB rate rises more rapidly over $z\leq 4$ than other star formation estimators Yüksel et al. (2008); Kistler et al. (2009); Ghirlanda & Salvaterra (2022). Furthermore, the local population of LGRB hosts ($z\leq 2$) tends towards lower mass hosts compared to the general population of star formatting galaxies (Krühler et al., 2015; Vergani et al., 2015; Japelj et al., 2016; Perley et al., 2016). With more complete and unbiased samples of LGRBs and their host galaxies, Greiner et al. (2015) have shown that between $z=3$ and $z=5$, the LGRB rate traces the UV metrics of the cosmic star formation. Together with average metallicity measurements of LGRB host galaxies, this has been interpreted as a low-metallicity preference for the formation of LGRBs (Levesque et al., 2010b, a; Graham & Fruchter, 2013; Perley et al., 2016; Vergani et al., 2017; Palmerio et al., 2019). A sharp cut-off of LGRB formation at a sub-solar metallicity can explain these observed phenomena over redshift, although its exact value ranges from $0.3\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ to $0.7\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ depending on the observational survey (Modjaz et al., 2008; Salvaterra et al., 2012; Graham & Fruchter, 2013; Perley et al., 2016; Vergani et al., 2017; Palmerio et al., 2019) and metallicity calibration used (Kewley & Ellison, 2008).
@@
Binary systems provide additional formation pathways for stars to gain angular momentum and be stripped during their evolution instead of requiring the star to be rotating at birth. On the other hand, processes such as common envelope evolution can remove angular momentum from the star and restrict the formation of LGRBs. In general, binary formation channels invoke chemically homogenous evolution or late-time tidal spin-up, although several other mechanisms have been proposed throughout the years (see Fryer et al., 2024, for an extensive overview). First, close massive binary systems can undergo mass transfer and spin up the accreting star sufficiently to undergo quasi-chemically homogeneous evolution at metallicities below $Z=1/3\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ (Cantiello et al., 2007; Yoon et al., 2008, 2010; Eldridge et al., 2011; Ghodla et al., 2023), as long as the accretor has not yet formed a well-defined stellar core. Secondly, in very tight near-equal-mass-ratio contact binaries, tides can spin up both stars, where the chemically homogenous evolution keeps both stars from expanding while forming the LGRB progenitors (Marchant et al., 2016; Song et al., 2016; Mandel & de Mink, 2016). Additionally, if the mass ratio is further from unity, just the initially more massive star in a binary can undergo this tidal spin up (de Mink et al., 2009; Marchant et al., 2017). Finally, tidal spin-up can also occur after a helium star has formed in tight binaries with a compact object companion (van den Heuvel & Yoon, 2007; Detmers et al., 2008; Qin et al., 2018). These four LGRB formation channels are more common at sub-solar metallicities.
@@ 2 Method
We use the public, state-of-the-art binary population synthesis code, POSYDON111 This work uses the POSYDON (posydon.org) version from commit 1e62bad. (Fragos et al., 2023; Andrews et al., 2024), which employs grids of detailed single- and binary-star models computed with the 1D stellar evolution code MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023) to create synthetic populations of evolved binary systems. POSYDON v2 release [^2] The POSYDON v2 grids will be available on the POSYDON Zenodo community. contains grids at eight metallicities: $10^{-4},10^{-3},10^{-2},0.2,0.45,0.1,1,2Z_{\odot}$ with $\mathrm{Z}_{\odot}=0.0142$, which include self-consistent modelling of internal rotation, as well as angular momentum transport in the stellar interior and between the stars and the orbit of the binary.
@@ 2.1 Stellar and binary evolution physics
The POSYDON H-rich+H-rich binary star grids, which start their evolution as two ZAMS, combine different formulations for mass transfer depending on the evolutionary state of the stars when Roche lobe overflow (RLOF) occurs (for details, see Fragos et al., 2023). For stars on the main sequence, the contact scheme is used, which keeps the stars inside their Roche lobe (Marchant et al., 2016). The kolb scheme (Kolb & Ritter, 1990) is used for stars undergoing RLOF post-main-sequence to allow for radial expansion past its Roche lobe since the envelope has a low density. The grids use the correction described in Xing et al. (2024) to allow mass transfer from the companion onto the primary star with the kolb scheme in MESA. Either star can initiate the mass transfer, and multiple mass-transfer phases can occur in the lifetime of a binary if the interaction remains stable. Stellar rotation enhances the stellar winds to keep rotation below the critical threshold $\omega_{s}/\omega_{\mathrm{s,crit}}\leq 1$. The latter is also a mechanism that can limit the accretion efficiency of a stable RLOF mass-transfer phase.
POSYDON considers mass transfer in the H-rich+H-rich grids as unstable either when (i) a donor mass transfer rate of $0.1\leavevmode\nobreak\ \mathrm{M}_{\odot}$ yr-1 is reached, or (ii) when mass is lost from the $L_{2}$ point, or (iii) when both stars fill their Roche lobe while either has evolved off the main sequence [^3] In the H-rich+H-rich grids, an additional unstable mass transfer condition is applied in later reruns of previously non-converged models. This condition only impacts low-mass accretors (see Section 4.1 in Andrews et al. (2024)) and falls outside the relevant mass range for this work. If a contact phase happens while both stars are on the main sequence, this is modelled following Marchant et al. (2016) and can result in a stable overcontact phase. MESA models reaching one of the unstable mass transfer criteria are stopped and evolved using POSYDON through a common envelope phase and subsequent detached or interaction phases. If none of the instability conditions are met, the binary evolution is tracked within MESA until the end of core carbon burning [^4] White dwarf and pair-instability progenitors are stopped earlier. See Fragos et al. (2023) and Andrews et al. (2024) for more details on the exact stopping criteria., where the final fate of each star as a neutron star or BH is determined. At this point, we determine whether it is an LGRB progenitor.
@@ 2.3 Binary population properties
We create stellar populations at each metallicity in POSYDON following a Kroupa (2001) initial mass functions between $7\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and $120\leavevmode\nobreak\ \mathrm{M}_{\odot}$, with a flat mass ratio, and log-period distributions. We initiate 1 000 00010000001\,000\,0001 000 000 binaries per metallicity and evolve them using the POSYDON nearest neighbour interpolation scheme.
@@ 3.1.1 Example model
Figure 1: A typical example of the evolution of a potential LGRB progenitor through the stable reverse-mass-transfer channel. The initial properties of the binary are $Z=\mathrm{Z}_{\odot}$, $M_{\mathrm{1,ZAMS}}\approx 25.1\leavevmode\nobreak\ \mathrm{M}_{\odot}$, $q=0.95$, and $P\approx 26.8$ days. In all panels, red (blue) solid lines refer to the primary (secondary) star. The columns are split into five different evolutionary phases, focussing on the mass transfer phases. The second and fourth columns show the mass transfer and ‘reverse’ mass transfer phases, respectively. The first, third and last columns show the long, nuclear timescale, detached evolution of the system. Top two rows: total mass (solid lines), He core mass (dashed lines) and CO core mass (dotted lines). The He (CO) core boundaries are set where the hydrogen (helium) fraction falls below 0.1. Third row: mass transfer rate from Roche lobe overflow. The periods of Roche lobe overflow are marked by a red (blue) shading when the donor is the primary (secondary). Fourth row: evolution of stellar radius (red: primary; blue: secondary) and separation of the system (black). Fifth row: surface rotational angular frequency in units of the critical (i.e. mass shed) value. Sixth row: the dimensionless spin parameter $cJ/GM^{2}$ of the star.
@@
Figure 3: Four grid slices at $\mathrm{Z}_{\odot}$ with $q=0.85$, $0.90$, $0.95$, and $0.99$. Orange symbols represent systems where only the primary fills its Roche lobe throughout the evolution of the binary model. Green squares indicate systems where also the secondary fills its Roche lobe, but not at the same time as the primary. Purple symbols indicate systems where both stars fill their Roche lobe at the same time. The type of symbol shows the stability of the interaction: squares indicate stable mass transfer, while binary models reaching instability criteria are indicated with a diamond. Grey squares indicate no mass transfer, while black dots indicate a non-converged model or Roche lobe overflow when the model is initiated. For systems where the companion fills its Roche lobe and the interaction remains stable, we indicate if an accretion disk is formed during the collapse of the model. This is indicated by a coloured circle inside the square symbol, where the colour indicates the size of the accretion disk.
@@ 3.1.2 Dependence on mass ratio, period and donor mass
In our example model, both stars fill their Roche lobe on the Hertzsprung gap after completing hydrogen burning in their cores. This moment of interaction is common in the proposed LGRB formation channel and requires the main-sequence lifetimes of the binary components to be similar. Otherwise, the primary star will have reached carbon exhaustion before the secondary has left the main sequence. This evolutionary timescale difference is influenced by the mass of the primary and the binary mass ratio. Specifically, the more equal the masses of the stars in the binary, the smaller the lifetime difference. As a result, many systems with $q$ close to unity have the secondary fill its Roche lobe before the primary completes its evolution. This effect can be seen in Figure 3, where an example grid slices of solar metallicity POSYDON detailed binary models with initial mass ratios of $q=0.85$, $q=0.90$, $q=0.95$, and $q=0.99$ are shown. The green markers indicate systems with the secondary expanding past its Roche lobe, while the pink markers indicate systems with only the primary star filling its Roche lobe. Systems with both stars filling their Roche lobe at the same time are marked with purple, as contact systems. As the mass ratio approaches $1$, the green area becomes larger indicating that more systems have reverse mass transfer. Additionally, the minimum mass ratio for the secondary to fill its Roche lobe depends on the absolute mass of the donor star. Massive stars are dominated by radiation pressure, which causes their main-sequence lifetimes to become similar towards higher masses. This means that binaries with more massive primaries have a broader range of mass ratios where the secondary can initiate a reverse mass transfer phase. As a result, the highest mass primaries in Figure 3, still have reverse mass transfer at $q=0.85$, while the lower mass systems do not.
@@
In Figure 3, it is important to address that the secondary does fill its Roche lobe in contact systems, but often reaches the $L_{2}$ overflow criteria for unstable mass transfer or, at wider periods, one of the stars leaves the main sequence while in contact. In general, only when the initial period is long enough can the secondary fill its Roche lobe without the binary reaching an unstable contact state.
@@
A stable mass transfer phase is a requirement to provide sufficient angular momentum to the primary star. In Figure 3 at $q=0.95$, two islands of unstable reverse mass transfer are present in the regime for Roche lobe overflow initiated by the secondary. The higher period island around $10^{3}$ days is a result of reaching the maximum mass transfer rate, while the lower period island near $10^{2}$ days is a result of $L_{2}$ overflow. While reverse mass transfer occurs more often nearing $q=1$, the grid at $=0.99$ in Figure 3 at solar metallicity shows that this is often unstable. Many binaries at $q=0.99$ reach a contact phase due to both stars evolving on a similar timescale, which causes many massive donors to reach contact off the main sequence even for large separations. The island of $L_{2}$ overflow remains and the diagonal unstable mass transfer is caused by the maximum mass transfer rate. There is still a large range of stable reverse mass transfer present at $q=0.99$, but not as extensive as at $q=0.95$ and many of the primary masses are too small to form a BH at collapse.
@@
The final limitation in the formation of LGRB through the stable reverse-mass-transfer channel is the angular momentum content of the primary at collapse. The circles inside the squares in grid slices in Figure 3 show the stable reverse mass transfer binaries that form an accretion disk during their collapse. Their colour indicates the size of the accretion disk that is formed. The $q=0.90$ and $q=0.95$ slices show a gradient in the disk mass with donor mass and period. Wider systems produce a smaller accretion disk than shorter-period binaries and the widest periods with stable reverse mass transfer do not produce a rapidly rotating star at collapse. This boundary moves to shorter periods as the donor mass increases.
In a long-period binary, the first mass transfer phase is shorter and leaves a larger hydrogen envelope on the primary than in a shorter-period system. As a result, the primary in the period-period system finds an equilibrium at a larger radius, which leads to higher angular momentum loss through the stellar winds compared to the less extended tight-binary primary. Once the reverse mass transfer phase is reached, the long-period primary has lost significantly more angular momentum. As the period becomes shorter, less hydrogen envelope remains after the initial mass transfer and, in most cases, no envelope is left, thus resulting in less angular momentum loss before the reverse mass transfer phase is reached.
Similarly to the first mass transfer, the reverse mass transfer has a longer duration in short-period binaries, which gives the primary time to partially stabilise and radially shrink. This allows for more mass and angular momentum accretion, which results in a higher total angular momentum in the primary of a short-period than a large-period binary. Although tides are stronger in the short-period binaries, they do not lead to sufficient angular momentum loss to prevent the formation of a disk at collapse. This restricts the LGRB formation in wide stable reverse-mass-transfer systems, which are also present in the $q=0.99$ grid slice.
@@
In two binaries with the same mass ratio and period but with different primary masses, the more massive primary will produce a more massive helium core. As a result, after the reverse mass transfer has taken place, it loses more mass and angular momentum due to stronger stellar winds compared to a less massive primary. A similar effect happens in wider binaries, where partial stripping can result in the formation of a slightly larger helium core and stronger mass loss after the reverse mass transfer phase. This effect is visible in Figure 3, where at the same period, an initially more massive star in the regime of stable reverse mass transfer produces a smaller or no accretion disk than a lower mass primary.

### v2

@@ 1 Introduction
Due to their highly-beamed and energetic nature, LGRBs have been observed up to $z\sim 9.4$ (Cucchiara et al., 2011) and have been used to estimate the star formation rate at high redshifts. However, the LGRB rate rises more rapidly over $z\leq 4$ than other star formation estimators (Yüksel et al., 2008; Kistler et al., 2009; Ghirlanda & Salvaterra, 2022). Furthermore, the local population of LGRB hosts ($z\leq 2$) tends towards lower mass hosts compared to the general population of star-forming galaxies (Krühler et al., 2015; Vergani et al., 2015; Japelj et al., 2016; Perley et al., 2016). With more complete and unbiased samples of LGRBs and their host galaxies, Greiner et al. (2015) have shown that between $z=3$ and $z=5$, the LGRB rate traces the UV metrics of the cosmic star formation. Together with average metallicity measurements of LGRB host galaxies, this has been interpreted as a low-metallicity preference for the formation of LGRBs (Levesque et al., 2010b, a; Graham & Fruchter, 2013; Perley et al., 2016; Vergani et al., 2017; Palmerio et al., 2019). A sharp cut-off of LGRB formation at a sub-solar metallicity can explain these observed phenomena over redshift, although its exact value ranges from $0.3\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ to $0.7\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ depending on the observational survey (Modjaz et al., 2008; Salvaterra et al., 2012; Graham & Fruchter, 2013; Perley et al., 2016; Vergani et al., 2017; Palmerio et al., 2019) and metallicity calibration used (Kewley & Ellison, 2008).
@@
Binary systems provide additional formation pathways for stars to gain angular momentum and be stripped during their evolution instead of requiring the star to be rotating at birth. On the other hand, processes such as common envelope evolution can remove angular momentum from the star and restrict the formation of LGRBs. In general, binary formation channels invoke chemically homogenous evolution or late-time tidal spin-up, although several other mechanisms have been proposed throughout the years (see Fryer et al., 2025, for an extensive overview). First, close massive binary systems can undergo mass transfer and spin up the accreting star sufficiently to undergo quasi-chemically homogeneous evolution at metallicities below $Z=1/3\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ (Cantiello et al., 2007; Yoon et al., 2008, 2010; Eldridge et al., 2011; Ghodla et al., 2023), as long as the accretor has not yet formed a well-defined stellar core. Secondly, in very tight near-equal-mass-ratio contact binaries, tides can spin up both stars, where the chemically homogenous evolution keeps both stars from expanding while forming the LGRB progenitors (Marchant et al., 2016; Song et al., 2016; Mandel & de Mink, 2016). Additionally, if the mass ratio is further from unity, just the initially more massive star in a binary can undergo this tidal spin up (de Mink et al., 2009; Marchant et al., 2017). Finally, tidal spin-up can also occur after a helium star has formed in tight binaries with a compact object companion (van den Heuvel & Yoon, 2007; Detmers et al., 2008; Qin et al., 2018). These four LGRB formation channels are more common at sub-solar metallicities.
@@ 2 Method
We use the public, state-of-the-art binary population synthesis code, POSYDON111 This work uses the POSYDON (posydon.org) version from commit 4a8c215. (Fragos et al., 2023; Andrews et al., 2025), which employs grids of detailed single- and binary-star models computed with the 1D stellar evolution code MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023) to create synthetic populations of evolved binary systems. POSYDON v2 release [^2] The POSYDON v2 grids will be available on the POSYDON Zenodo community. contains grids at eight metallicities: $10^{-4},10^{-3},10^{-2},0.2,0.45,0.1,1,2Z_{\odot}$ with $\mathrm{Z}_{\odot}=0.0142$, which include self-consistent modelling of internal rotation, as well as angular momentum transport in the stellar interior and between the stars and the orbit of the binary.
@@ 2.1 Stellar and binary evolution physics
The POSYDON H-rich+H-rich binary star grids, which start their evolution as two ZAMS stars, combine different formulations for mass transfer depending on the evolutionary state of the stars when Roche lobe overflow (RLOF) occurs (for details, see Fragos et al., 2023). This grid self-consistently treats rotation, tides, and mass transfer within the binary, as the orbital evolution and stellar structure equations of both stars are solved with MESA. For stars on the main sequence, the contact scheme is used, which keeps the stars inside their Roche lobe (Marchant et al., 2016). The kolb scheme (Kolb & Ritter, 1990) is used for stars undergoing RLOF post-main-sequence to allow for radial expansion past its Roche lobe since the envelope has a low density. The grids use the correction described in Xing et al. (2024) to allow mass transfer from the companion onto the primary star with the kolb scheme in MESA. Either star can initiate the mass transfer, and multiple mass-transfer phases can occur in the lifetime of a binary if the interaction remains stable. Stellar rotation enhances the stellar winds to keep rotation below the critical threshold $\omega_{s}/\omega_{\mathrm{s,crit}}\leq 1$. The latter is also a mechanism that can limit the accretion efficiency of a stable RLOF mass-transfer phase.
POSYDON considers mass transfer in the H-rich+H-rich grids as unstable either when (i) a donor mass transfer rate of $0.1\leavevmode\nobreak\ \mathrm{M}_{\odot}$ yr-1 is reached, or (ii) when mass is lost from the $L_{2}$ point, or (iii) when both stars fill their Roche lobe while either has evolved off the main sequence [^3] In the H-rich+H-rich grids, an additional unstable mass transfer condition is applied in later reruns of previously non-converged models. This condition only impacts low-mass accretors (see Section 4.1 in Andrews et al. (2025)) and falls outside the relevant mass range for this work. If a contact phase happens while both stars are on the main sequence, this is modelled following Marchant et al. (2016) and can result in a stable overcontact phase. Binary MESA models reaching one of the unstable mass transfer criteria are stopped and evolved using POSYDON through a common envelope phase and subsequent detached or interaction phases. If none of the instability conditions are met, the binary evolution is tracked within MESA until the end of core carbon burning [^4] White dwarf and pair-instability progenitors are stopped earlier. See Fragos et al. (2023) and Andrews et al. (2025) for more details on the exact stopping criteria., where the final fate of each star as a neutron star or BH is determined. At this point, we determine whether it is an LGRB progenitor.
@@ 2.3 Binary population properties
We create stellar populations at each metallicity in POSYDON following a Kroupa (2001) initial mass function between $7\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and $120\leavevmode\nobreak\ \mathrm{M}_{\odot}$, with a flat mass ratio, and log-period distributions. We initiate 1 000 00010000001\,000\,0001 000 000 binaries per metallicity and evolve them using the POSYDON nearest neighbour interpolation scheme.
@@ 3.1.1 Example model
Figure 1: A typical example of the evolution of a potential LGRB progenitor through the stable reverse-mass-transfer channel. The initial properties of the binary are $Z=\mathrm{Z}_{\odot}$, $M_{\mathrm{1,ZAMS}}\approx 25.1\leavevmode\nobreak\ \mathrm{M}_{\odot}$, $q=0.95$, and $P\approx 26.8$ days. In all panels, red (blue) solid lines refer to the primary (secondary) star. The columns are split into five different evolutionary phases, focusing on the mass transfer phases. The second and fourth columns show the mass transfer and ‘reverse’ mass transfer phases, respectively. The first, third and last columns show the long, nuclear timescale, detached evolution of the system. Top two rows: total mass (solid lines), He core mass (dashed lines) and CO core mass (dotted lines). The He (CO) core boundaries are set where the hydrogen (helium) fraction falls below 0.1. Third row: mass transfer rate from Roche lobe overflow. The periods of Roche lobe overflow are marked by a red (blue) shading when the donor is the primary (secondary). Fourth row: evolution of stellar radius (red: primary; blue: secondary) and separation of the system (black). Fifth row: surface rotational angular frequency in units of the critical (i.e. mass shed) value. Sixth row: the dimensionless spin parameter $cJ/GM^{2}$ of the star.
@@
Figure 3: Four grid slices at $\mathrm{Z}_{\odot}$ with $q=0.85$, $0.90$, $0.95$, and $0.99$. Orange symbols represent systems where only the primary fills its Roche lobe throughout the evolution of the binary model. Green squares indicate systems where also the secondary fills its Roche lobe, but not at the same time as the primary. Purple symbols indicate systems where both stars fill their Roche lobe at the same time. The type of symbol shows the stability of the interaction: squares indicate stable mass transfer, while binary models reaching instability criteria are indicated with a diamond. Grey squares indicate no mass transfer, while black dots indicate a non-converged model or Roche lobe overflow when the model is initiated. For systems where the companion fills its Roche lobe and the interaction remains stable, we indicate if an accretion disk is formed during the collapse of the model. This is indicated by a coloured circle inside the square symbol, where the colour indicates the mass of the accretion disk.
@@ 3.1.2 Dependence on mass ratio, period and donor mass
In our example model, both stars fill their Roche lobe on the Hertzsprung gap after completing hydrogen burning in their cores. This moment of interaction is common in the proposed LGRB formation channel and requires the main-sequence lifetimes of the binary components to be similar. Otherwise, the primary star will have reached carbon exhaustion before the secondary has left the main sequence. This evolutionary timescale difference is influenced by the mass of the primary and the binary mass ratio. Specifically, the more equal the masses of the stars in the binary, the smaller the lifetime difference. As a result, many systems with $q$ close to unity have the secondary fill its Roche lobe before the primary completes its evolution. This effect can be seen in Figure 3, where example grid slices of solar metallicity POSYDON detailed binary models with initial mass ratios of $q=0.85$, $q=0.90$, $q=0.95$, and $q=0.99$ are shown. The green markers indicate systems with the secondary expanding past its Roche lobe, while the orange markers indicate systems with only the primary star filling its Roche lobe. Systems with both stars filling their Roche lobe at the same time are marked with purple, as contact systems. As the mass ratio approaches $1$, the green area becomes larger, indicating that more systems have reverse mass transfer. Additionally, the minimum mass ratio for the secondary to fill its Roche lobe depends on the absolute mass of the donor star. Massive stars are dominated by radiation pressure, which causes their main-sequence lifetimes to become similar towards higher masses. This means that binaries with more massive primaries have a broader range of mass ratios where the secondary can initiate a reverse mass transfer phase. As a result, the highest mass primaries in Figure 3, still have reverse mass transfer at $q=0.85$, while the lower mass systems do not.
@@
In Figure 3, it is important to address that the secondary does fill its Roche lobe in contact systems, but often reaches the $L_{2}$ overflow criteria for unstable mass transfer or, at longer periods, one of the stars leaves the main sequence while in contact. In general, only when the initial period is long enough can the secondary fill its Roche lobe without the binary reaching an unstable contact state.
@@
A stable mass transfer phase is a requirement to provide sufficient angular momentum to the primary star to produce an accretion disk at collapse. In Figure 3 at $q=0.95$, two islands of unstable reverse mass transfer are present in the regime for Roche lobe overflow initiated by the secondary. The higher period island around $10^{3}$ days is a result of reaching the maximum mass transfer rate, while the lower period island near $10^{2}$ days is a result of $L_{2}$ overflow. While reverse mass transfer occurs more often nearing $q=1$, the grid at $q=0.99$ in Figure 3 at solar metallicity shows that this is often unstable. Many binaries at $q=0.99$ reach a contact phase due to both stars evolving on a similar timescale, which causes many massive donors to reach contact off the main sequence even for large separations. The island of $L_{2}$ overflow remains and the diagonal unstable mass transfer is caused by the maximum mass transfer rate. There is still a large range of stable reverse mass transfer present at $q=0.99$, but not as extensive as at $q=0.95$ and many of the primary masses are too small to form a BH at collapse.
@@
The final limitation in the formation of LGRB through the stable reverse-mass-transfer channel is the angular momentum content of the primary at collapse. The circles inside the squares in grid slices in Figure 3 show the stable reverse-mass-transfer binaries that form an accretion disk during their collapse. Their colour indicates the mass of the accretion disk that is formed. The $q=0.90$ and $q=0.95$ slices show a gradient in the disk mass with donor mass and period. In general, wider systems produce a less massive accretion disk than shorter-period binaries, and the widest periods with stable reverse mass transfer do not produce a rapidly rotating star at collapse. The boundary of accretion disk formation moves to shorter periods as the donor mass increases and is a result of hydrogen being present when the reverse mass transfer occurs, which inhibits the transfer of angular momentum into the core of the primary.
We explore the gradient in the accretion disk mass as a function of donor mass and period in Appendix A. In short, at the same mass and mass ratio, the accreted angular momentum and mass during the stable reverse-mass-transfer phase decrease with increasing period due to the mass transfer being shorter and, thus, there being less time to diffuse the angular momentum into deeper layers of the accretor. After the stable reverse-mass-transfer phase, the fraction of angular momentum loss is similar across different initial periods, around 80%. Thus, short-period systems, which gained more angular momentum through mass transfer, reach core carbon depletion with a larger angular momentum reservoir, sufficient to produce an accretion disk during collapse.
@@
The gradient as a function of donor mass originates from a slightly different physical phenomenon. At the same period and mass ratio with an increasing initial donor mass, more angular momentum is gained by the primary during the reverse-mass-transfer phase. However, the initially more massive primary loses more angular momentum and mass due to stellar winds after the stable reverse mass transfer. A lower total angular momentum amount and the increased total mass at core carbon depletion limit the mass of the formed accretion disk. This effect is visible in Figure 3, where at the same period, an initially more massive star in the regime of stable reverse mass transfer produces a smaller or no accretion disk compared to a lower mass primary.